OpenStudy (anonymous):

a line with slope 3 pasing throught the point(4,-2) find the slope intercept form how do u work this out

OpenStudy (anonymous):

if slope = 3, then it means m = 3 y = mx+b ,where m = 3 now just substitute the given point: (4,-2) -2 = 3*(4) +b b = -14 equation: y = 3x-14

OpenStudy (anonymous):

dam thanks